OpenStudy (anonymous):

Choose the point on the terminal side of -210°. I have gotten down to the reference angle 180-150 = 30 and then I don't know what to do from there.

OpenStudy (anonymous):

\[cos(-210), sin(-210)=(-\frac{\sqrt{3}}{2},-\frac{1}{2})\]